OpenStudy (anonymous):

Jack and Jill wanted to meet in Lansing to watch the Mi state basketball game. Jack was comin from flint and jill from grand rapids. Jaxk drove at an averad speed of 6o mph while Jill drove at an avrage speed at 70 mph. How long will it to take Jack get to lansin if he is 40 miles away?

OpenStudy (anonymous):

40miles/60mph = 2/3 hours?

OpenStudy (anonymous):

right that's what I thought but 40/60 is not 2/3 of n hour

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

confuseed? 1 seeec.

OpenStudy (anonymous):

40/60 = 2/3

OpenStudy (anonymous):

I got 0.6666666

OpenStudy (anonymous):

O dummbyy(me) Im reduciinng Im sooo dummb thankkss botthh :)

OpenStudy (anonymous):

=D

OpenStudy (anonymous):

np

OpenStudy (anonymous):

1 more questiion How long would it take jill to get to lasin if she is 105 miles away

OpenStudy (anonymous):

105/70

OpenStudy (anonymous):

Needhelps, do you know how dividing units work? It'll be really helpful for these types of problems. E.g. miles/ (miles/hour) = hour

OpenStudy (anonymous):

1 hour 30 min?

OpenStudy (anonymous):

I queess but I'm pretty solid on dividing I'm just confused..

OpenStudy (anonymous):

ah, that's why you should check units!

OpenStudy (anonymous):

okkay/